Breakfasttime Cereal Company manufactures two breakfast cereals in a joint process. Cost and quantity information is as follows:

    

Joint Cost Cereal Quantity at Split-Off Point Sales Price per Kilogram
$ 34,000 Yummies 18,900 kilograms $ 3.00
Crummies 8,100 kilograms 3.50

   

Required:

Use the physical-units method to allocate the company’s joint production cost between Yummies and Crummies. (Do not round intermediate calculations.)

Allocation of Joint Cost
Yummies
Crummies
Answer #1

Total physical units = 18,900 + 8,100 = 27,000

Allocation of joint cost :-

Yummies = 34,000 x (18,900/27,000) = $23,800

Crummies = 34,000 x (8,100/27,000) = $10,200